OpenStudy (anonymous):

Given ycos(x^2) = xcos(y^2), use implicit differentiation to find dy/dx.

OpenStudy (anonymous):

phewww i hope this right ;)

OpenStudy (raden):

at the 1st steps, u were mistake... so on also be mistake :) look at the left side, it should be y'cos(x^2) - 2xysin(x^2)

OpenStudy (anonymous):

you right

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

so the rest is screwed off lol

OpenStudy (raden):

when differentiate of y, so cos(x^2) will be coeff, doesnt it

OpenStudy (raden):

lol, yes

OpenStudy (anonymous):

but the right side I think i got it?

OpenStudy (raden):

yeah, for right side very good i thnk..

OpenStudy (anonymous):

lol

OpenStudy (raden):

np, sometimes i have done mistake too like u :)

OpenStudy (anonymous):

\[\frac{ \cos(y^2) +2xysin(x^2)}{ \cos(x^2)+2xysin(y^2)}\] is this what you get?

OpenStudy (raden):

i got dy/dx = (2xysin(x^2) + cos(y^2)) / (2xysin(y^2) + cos(x^2)) if i dont mistake also :)

OpenStudy (anonymous):

yeeh then we got it ;;)..long arse problem (*^_~*)

OpenStudy (raden):

hoho... but i think asker be angry if we just discuss the answer only without show the steps for solution :D

OpenStudy (anonymous):

\[y'\cos(x^2)-ysin(x^2)(2x)=\cos(y^2)+x(-\sin(y^2))(2y)y'\] \[y'\cos(x^2)-2xysin(x^2)=\cos(y^2)-2xysin(y^2)y'\] \[y'\cos(x^2)+2xysin(y^2)y'=\cos(y^2)+2xysin(x^2)\] \[y'(\cos(x^2)+2xysin(y^2))=\cos(y^2)+2xysin(x^2)\] \[y'=\frac{ \cos(y^2)+2xysin(x^2) }{ \cos(x^2)+2xysin(y^2) }\]